OpenStudy (anonymous):

Compare and Contrast. Below are two sets of points. Find the slopes between the two points and compare them. Choose the statement that is true about the values of the slopes. Set #1 Set #2 (2, 4) and (1, 1) (4, 6) and (2, 7) The slope of Set #1 is smaller than the slope of Set #2. The slope of Set #1 is larger than the slope of Set #2. The slope of Set #1 is the same as the slope of Set #2. None of the statements above describes slopes of sets shown.

OpenStudy (anonymous):

Do you know how to get the slope between a pair of points?

OpenStudy (anonymous):

\[\frac{ y _{1}- y _{2} }{ x _{1} - x _{2} } = m\]and "m" is the slope. You use points (x1, y1) and (x2, y2) for your values. Substitute and simplify. So same for the other pair of points.

OpenStudy (anonymous):

Is this making sense to you?

OpenStudy (anonymous):

a little

OpenStudy (anonymous):

isn it A

OpenStudy (anonymous):

ok, that's a start your points (2, 4) and (1, 1) are in the form (x, y) with (2, 4) being (x1, y1) and (1, 1) being (x2, y2). So, x1=2, y1=4, x2=1, and y2=1. Study how those values where obtained. Then you can make the substitutions into the equation from my first post.

OpenStudy (anonymous):

ion think none of the statements above decribe the slopes

OpenStudy (anonymous):

So, if you use the points and values from my last post, you would be able to get the slope of the first pair of points. What did you get for that slope value?

OpenStudy (anonymous):

im still confused i have a hard time understanding these things....sorry

OpenStudy (anonymous):

Try writing the equation out here in this thread. You don't have to simplify the equation yet. We'll take it slow. Just do the substitutions for now and then we can get the slope. Show me what you've got so far.

OpenStudy (anonymous):

i dont knw non of this stuff

OpenStudy (anonymous):

We'll get you there, don't worry. In my fourth post I wrote that y1=4 and y2=1. Using those variables, can you write an expression for y1 - y2, making the substitutions?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

That's the spirit! Ok, now, Write that out here and you'll be amazed at what you'll be able to do.

OpenStudy (anonymous):

y2,y1 M=----- x2,x1

OpenStudy (anonymous):

2 and -2 than i divided and got -1 just like the last.. so it is b right?

OpenStudy (anonymous):

We'll back up a little more and then you'll get it. Sometimes it requires just getting over math anxiety. If I tell you y1 = 4, then I'm saying that y1 is the name of the variable and its value is 4. If I ask you to tell me what the value of y1 is, can you do this?

OpenStudy (anonymous):

-1 right?

OpenStudy (anonymous):

I know you want to be able to get to the answer, and I'm going to try to get you there, believe me. But you will have to go through a couple steps first. I know no one wants to just ask questions and get answers. You really want to be able to do this stuff on your own. y1 = 4. What is the value of y1? What does it equal?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Good! Believe me when I tell you this, that you just made progress. Some people struggle with the concept of a variable. It's a box with something inside of it. y1 box has a 4 inside of it. Now, y1=4, and y2=1 . What is the expression for y1 - y2? Just substitute the numbers.

OpenStudy (anonymous):

jus add the numbers

OpenStudy (anonymous):

Well, there is a "-" sign between the numbers and "-" means that you have to subtract, so the second number gets subtracted from the first one.

OpenStudy (anonymous):

0

OpenStudy (anonymous):

What is y2?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

=y2-y/x2-x =y(y-1)/x(x-1) =y(x-1)x(y-1) =(xy-y)(xy-x) =x2y2-x2y-xy2+xy =>xy(xy-x-y+1)

OpenStudy (anonymous):

It's probably not fair to you for us to continue. I want to help you, but I'm not convinced that you are getting much out of online tutoring, though I admire your effort and willingness to learn. Perhaps you might benefit from live tutoring. I encourage you to find a good live tutor in your area. Good luck to you and I sincerely hope you eventually get this. Take care .